• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Kalyan

    Log entry time 14:11:44 on February 26, 2011

    Entry number 348826

    keyword=Hall status

    Techs are still working on fixing the spectrometer bogies while MCC is working on first pass spilt for both Hall-A and C.